OpenStudy (anonymous):

factor the expression completely 12^2+36+27

OpenStudy (anonymous):

something missing?

OpenStudy (anonymous):

12x^2+36x+27

OpenStudy (soni):

144x+36x+27= 144x+36x=180x 180x /27 =6.6